OpenStudy (anonymous):

Find the length of the radius for this circle. (x - 3)² + (y - 5)² = 8

OpenStudy (anonymous):

Ok, so this is already in the form of a circle. The generic form of a circle is: \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

In your case, \[r^2 = 8\]

OpenStudy (anonymous):

So, solving for r, you get \[\pm \sqrt{8}\]

OpenStudy (anonymous):

Of course, it makes no sense to have a negative radius, so you take the positive answer, \[r=\sqrt{8}\]